OpenStudy (pandateddygirl):

Which equation is graphed here? A: y + 5 = -3/2 (x-4) B: y - 5 = -2/3 (x + 4) C: y - 5 = -3/2(x + 4) D: y - 4 = -3/2 ( x + 5)

OpenStudy (anonymous):

you got the slope?

OpenStudy (pandateddygirl):

no.....

OpenStudy (anonymous):

you gotta get that first

OpenStudy (anonymous):

you can do it with your eyeballs i can try to explain how

OpenStudy (pandateddygirl):

ok

OpenStudy (anonymous):

oh pick a point on the line that is on a corner doesn't matter which one you pick i pick \((0,1)\)

OpenStudy (pandateddygirl):

-3/2

OpenStudy (anonymous):

then see how you get to the next one that is on a corner you go to the right two, down 3 to get to the next one at \((2,-2)\)

OpenStudy (anonymous):

right 2, down 3, slope is \[-\frac{3}{2}\]

OpenStudy (anonymous):

Or you could choose two random points and use the formula. \[m=\frac{ y2-y1 }{ x2-x1} \]

OpenStudy (anonymous):

so now we know it is A, C or D didn't eliminate much

OpenStudy (pandateddygirl):

lol

OpenStudy (anonymous):

if it is y + 5 = -3/2 (x-4) that means the point \((4,-5)\) is on the graph is it?

OpenStudy (anonymous):

and the answer is no, it is not, so now we got rid of A

OpenStudy (anonymous):

if it is y - 5 = -3/2(x + 4) that means the point \((-4,5)\) is on the graph is it?

OpenStudy (pandateddygirl):

no

OpenStudy (anonymous):

then i hope it is the last one

OpenStudy (anonymous):

y - 4 = -3/2 ( x + 5) means the point \((-5,4)\) is on the graph i sure hope so!

OpenStudy (anonymous):

actually it is not did i make some mistake?

OpenStudy (anonymous):

ah yes i did \[(4,-5)\] IS on the graph

OpenStudy (anonymous):

i said "no" but it should have been "yes" go with A, i will get my eyes checked

OpenStudy (pandateddygirl):

lol thanks!